OpenStudy (anonymous):

Eight children from the Schickel family are combining their savings to purchase a gaming system to connect to and play on their television. The gaming system is on sale at their favorite discount electronics store for 10 percent off of its regular price of $500. If the cost of the discounted gaming system, excluding tax, is split equally among the eight children, how much money is each child to contribute? A $10.00 B $56.25 C $62.50 D $450.00

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

B

OpenStudy (anonymous):

explain it

OpenStudy (mathstudent55):

First, let's find the amount of the discount. 10% of 500 means 10% * 500. % means "divided by 100", so 10% means 10/100 = 1/10 = 0.1 That means that 10% of 500 means 0.1 * 500. Can you multiply to find how much the discount is?

OpenStudy (anonymous):

because 10/100 of 500 is 50 and 500 - 50 = 450. 450/8 =56.25

OpenStudy (anonymous):

need more elaboration?

OpenStudy (anonymous):

Ali drinks three 12-ounce sodas every day and has done so for the past 2 years. His dentist warns him that acid and sugar from drinking that many sodas are damaging his teeth. Ali needs to reduce his intake to two 12-ounce sodas per week. By what percent does Ali need to reduce his soda intake for the week? A 90% B 67% C 45% D 80%

OpenStudy (mathstudent55):

Let's look at how much he has been drinking each week for the past 2 years, and how much he should be drinking each week from now on. He drinks 3 sodas a day. How many sodas does he drink in one week?

OpenStudy (anonymous):

so

OpenStudy (anonymous):

.....

OpenStudy (mathstudent55):

3 sodas a day for a week is how many sodas in one week?

OpenStudy (anonymous):

21

OpenStudy (mathstudent55):

Great. He currently drinks 21 sodas per week. He needs to drink only 2 sodas per week. We need to find the percent change from 21 to 2.

OpenStudy (mathstudent55):

Percent change formula: \( percent ~change = \dfrac{new ~number - old ~number}{old ~number} \times 100\) A negative percent change is a percent decrease. A positive percent change is a percent increase. Use 2 for "new number" and 21 for "old number" in the formula above.

OpenStudy (anonymous):

5 Corbin is adding supporting beams to the legs of a table. If angle x and angle y are complementary angles, and angle x measures 17 degrees, what is the measurement of angle y? A 17 degrees B 163 degrees C 73 degrees D 90 degrees 6 Jennifer works in the child care center at her church. Children at the child care center play with building blocks. The blocks are kept in a container that has 2 congruent circular bases. The container is in the shape of a - A cone B triangular pyramid C cylinder D rectangular prism Page 3 Go On2013-2014 7B_U4L6 Additional Problems MA:07 7 Rehka’s father spends $50 on materials needed for a father-daughter project. He and his daughter build a bird house for the backyard, as shown. Which value best represents the amount of space a family of birds living in the house would have? A 352 cubic inches B 512 cubic inches C 608 cubic inches D 704 cubic inches

OpenStudy (anonymous):

have an answer